A kid spends 5/18 of his sum of money and 10$.What sum of money did this kid have nowing that 2/3 of his sum stays with him ?
I really need help !

Answer by stanbon(75887)

A kid spends 5/18 of his sum of money and 10$.What sum of money did this kid have nowing that 2/3 of his sum stays with him ?
---------------
Let amount of money starts with be "x" dollars.
-----------
EQUATION:
x - (5/18)x-10 = (2/3)x
(13/18)x - 10 = (12/18)x
13x - 180= 12x
x = $180 (Amount of money he started with).
======================
Cheers,
Stan H.